8Consider first the goods market model with constant investment that we saw in
Chapter 3. Consumption is given by:
C = Co + c1(Y-T)
And I, G and T are given.
a. Solve for equilibrium output. What is the value of the multiplier?
Now let investment depend on both sales and the interest rate:
I = b0 + b1Y -b2i
b. Solve for equilibrium output using the methods learned in chapter 3. At a
given interest rate, why is the effect of a change in autonomous spending
bigger than what it was in part (a)? Why? (Assume c1 + b1 = 1)
c.Solve for equilibrium level of investment.
d. Let’s go behind the scene in the monetary market. Use the equilibrium in the
money market M/P = d1Y – d2i to solve for the equilibrium level of the real
money supply.
How does the real money supply vary with government spending?
1
Expert's answer
2020-04-29T09:27:53-0400

a. The equilibrium output is:

"Y = C + G + I = Co + c1(Y-T) + G + I"

The value of the multiplier is: "m = \frac{1}{1 - c1}."

I = b0 + b1Y -b2i

b. The equilibrium output is:

"Y = C + G + I = Co + c1(Y-T) + G + b0 + b1Y - b2i".

At a given interest rate, the effect of a change in autonomous spending is bigger than what it was in part (a) because the multiplier effect now is higher.

c. In equilibrium investment equals saving, so:

"I = S,"

"b0 + b1Y - b2i = S."

d. In equilibrium in the money market money demand equals money supply, so:

"M\/P = Ms = d1Y \u2013 d2i."

The real money supply increases if government spending increases.